OpenStudy (anonymous):

What is the equation of the line that passes through the point (4, 1) and has a slope of 5?

OpenStudy (anonymous):

equation of the line is y-y1 = m(x-x1) where m = slope of the line and (x1,y1) = point through which the line passes

OpenStudy (anonymous):

y=5x-19

OpenStudy (anonymous):

Thank you!